# Equivariant Flow Matching for Symmetry-Breaking Bifurcation Problems

### v2

@@ Abstract
Bifurcation phenomena in nonlinear dynamical systems often lead to multiple coexisting stable solutions, particularly in the presence of symmetry breaking. Deterministic machine learning models are unable to capture this multiplicity, averaging over solutions and failing to represent lower-symmetry outcomes. In this work, we formalize the use of generative AI, specifically flow matching, as a principled way to model the full probability distribution over bifurcation outcomes. Our approach builds on existing techniques by combining flow matching with equivariant architectures and an optimal-transport-based coupling mechanism. We generalize equivariant flow matching to a symmetric coupling strategy that aligns predicted and target outputs under group actions, allowing accurate learning in equivariant settings. We validate our approach on a range of systems, from simple conceptual systems to physical problems such as buckling beams and the Allen-Cahn equation. The results demonstrate that the approach accurately captures multimodal distributions and symmetry-breaking bifurcations. Moreover, our results demonstrate that flow matching significantly outperforms non-probabilistic and variational methods. This offers a principled and scalable solution for modeling multistability in high-dimensional systems.
@@ 1 Introduction
Machine learning enables efficient data-driven modeling of complex systems, especially when traditional methods impose high complexity or are incomplete senior2020improved, brunton2020machine, pfaff2020learning, hendriks2025similarity, minartz2025discovering, brinke2025flow. Many such systems exhibit bifurcations that mark sudden changes in system behavior due to a small change in a control parameter. Bifurcations are crucial in fields such as fluid dynamics ruelle1971nature, climate science armstrong2022exceeding, willcock2023earlier, rietkerk2021evasion, steffen2018trajectories, biology sadria2024fatenet, and crowd dynamics muramatsu1999jamming, gu2025emergence, warren2022bifurcation, where predicting transitions is essential. Many bifurcating systems exhibit multistability, where multiple distinct stable states coexist under the same input parameters, and the number of these states can change as the parameters vary. This phenomenon often arises due to symmetry breaking, where a system’s state after the bifurcation possesses fewer symmetries than before. This results in a multiplicity of possible solutions, all of which are equally valid.
@@
Simulating such systems with machine learning is nontrivial and remains an open problem [zhang2025artificial, §2.9.1]. Deterministic machine learning models can only produce the average of multiple solutions, resulting in nonphysical predictions that do not represent any true solution. Geometric deep learning approaches that rely on equivariant models preserve the system’s symmetry but cannot select asymmetric outcomes, limiting their ability to capture bifurcations. Approaches such as canonicalization lawrence2025improving, input perturbation hendriks2025similarity, model ensembles zou2025learning, deterministic machine learning models combined with pseudo-arclength continuation fabiani2025enabling, della2025surrogate, and sampling group operations to artificially break symmetry xie2024equivariant, lawrence2025improving offer partial solutions but often lack generalizability or precision.
@@
We propose modeling the full output distribution using generative modeling. Generative models aim to learn mappings from simple base distributions (e.g., Gaussian noise) to complex, high-dimensional target distributions. However, the probability distributions we require are singular: the set of allowed values (the support) lies on a subspace that has lower dimensionality than the input space, e.g., a Dirac delta in 1D, or a circle in 2D space. This means the probability mass must be concentrated as sharply as possible around the allowed values. When the target distribution is singular, the necessary mapping becomes highly nonlinear: nearby points in the source distribution may map to distant points in the target space. Capturing such mappings is challenging for neural networks, as it requires representing high-frequency functions rahaman2019spectral. This is a fundamental limitation of direct generative approaches, which must learn a single, highly nonlinear transformation. For example, variational autoencoders (VAE) tend to generate blurry images due to their limitations in capturing highly concentrated, multimodal distributions larsen2016autoencoding, bousquet2017optimal, dosovitskiy2016generating.
@@
In contrast, iterative methods such as diffusion models, denoising processes, and flow matching approximate this complex mapping as a sequence of small integration steps. This iterative structure distributes the required nonlinearity across many stages, making the learning problem more tractable, as each step only needs to model a small, smooth transformation. As a result, these methods are better suited for (i) modeling singular distributions, where the probability mass is concentrated on a low-dimensional manifold, and (ii) multimodal distributions, where the probability mass must split far apart salmona2022can.
@@
Therefore, to address the challenge of modeling multistability, we utilize flow matching lipman2022flow to model the full output distribution. We show that flow matching excels at parameterizing distributions whose probability mass is highly concentrated in a low-dimensional subspace (e.g., very close to two Dirac deltas). We further systematically integrate flow matching with equivariant modeling to tackle symmetry-breaking. Additionally, we develop ‘symmetric coupling’, an approach to determine the optimal training target, similar to minimatch optimal transport tong2023improving and equivariant flow matching klein2023equivariant, song2023equivariant. This approach substantially improves the training and the quality of generated flow paths. We demonstrate our method on abstract and physical systems, including coin flips, buckling beams, and phase separation via the Allen-Cahn equation. Our approach effectively captures bifurcations and multistability, with and without symmetry breaking.
@@ 2 Methods
If an input $x$ is invariant (i.e., self-similar) under a subset of the group actions (i.e., $g_{x}\cdot x=x$ for all $g_{x}\in G_{x}$, with $G_{x}\subseteq G$), then equivariance implies that output $y$ is also invariant (i.e., $g_{x}\cdot y=y$). However, this preservation of symmetry inherently prevents a model from expressing symmetry-breaking behavior. To capture bifurcations and transitions into asymmetric states, equivariant models must be adapted to allow for symmetry-breaking solutions.
@@
The equivariance in Equation (1) then still holds for the set of solutions $\{g_{x}\cdot y|g_{x}\in G_{x}\}$.
This is similar to the definition of ‘relaxed equivariance’ by Kaba et al. kaba2023equivariance, kaba2023symmetry in that both account for input symmetries via stabilizer subgroups (i.e., groups $G_{x}$ whose actions leave the input $x$ invariant), but it differs by preserving equivariance over sets of outputs (i.e., orbits of solutions), whereas relaxed equivariance modifies the equivariance condition itself, by allowing a correction with an element from the input’s stabilizer subgroup.
As a probability distribution:
@@
To ensure that the predictions by the model satisfy Equation (2), we use (i) a model that is equivariant under $G$ to parametrize the probability distribution and (ii) a prior $p(y_{0})$ that is invariant under $G$. This ensures that the generated probability distribution respects the symmetry of the problem minartz2023equivariant.
@@
To model this probability distribution, we use flow matching lipman2022flow, which learns a vector field $u(y_{t},t,x)$ that transforms samples from an uninformed prior $p(y_{0})$ into samples $y_{1}$ from the target distribution $p(y|x)$ over a continuous pseudo-time variable $t\in[0,1]$. The vector field is learned by minimizing the expected squared error between the model prediction and the target vector field along linear interpolation paths between prior and target samples.
Symmetric coupling:
@@
This is a way of straightening the flow paths similar to minibatch optimal transport tong2023improving, but instead of finding the optimal coupling between a minibatch of the prior samples and target samples, we find the optimal coupling between one prior sample $y_{0}$ and all symmetric equivalents $g_{x}\cdot y_{1},g_{x}\in G_{x}$ of its corresponding target sample $y_{1}$. Our approach is similar to equivariant flow matching as described by Klein klein2023equivariant and Song song2023equivariant, except we only consider the group $G_{x}$ that leaves the input $x$ invariant. If this group contains permutations or rotations, we can therefore follow their lead and use the Hungarian algorithm for permutations or the Kabsch algorithm for rotations. Other options are checking all possible reflections of the target or using FFT cross-correlation to determine the optimal periodic translation of the target.
@@
Figure 1: (a) Self-similarity and equivariance illustrated with the buckling beam problem, in which the input is the state at the current time step $t$ and the target output is the next time step $t+1$. $d$ is the relative vertical downward displacement of the beam tip, which is increased over time. The relevant symmetry group $G$ here consists of only two elements: identity and reflection in the y-axis. In this case, the self-similarity of the beam before buckling happens to be described by the same group, i.e., here $G_{x}=G$. (b) The $x$-coordinates of each node over the course of a trajectory, showing 50 predictions of a trained flow matching model (dashed lines) compared to the two possible ground truths (solid lines). (c) The same 50 predictions compared to the two possible ground truths, showing the beam deformation.
@@ 3 Experiments
For the 4-node graph, we additionally show how symmetric coupling further improves the results. Symmetric coupling in this case amounts to picking the permutation of the target that has the smallest Euclidean distance squared to the noise sample. Only two permutations are relevant, because there are only two possible solutions; the identity permutation and the one that swaps node 0 with 2 and node 1 with 3.
Table 1: Performance of various approaches for the different test systems. All metrics except for the Allen-Cahn are the Wasserstein distance of a set of 100 predictions to the actual distribution of allowed outcomes for each test data point, averaged over all test data points. For the Allen–Cahn system, we use the residual of the governing equation as a performance metric, acknowledging its limitations due to sensitivity to small fluctuations in the prediction, particularly in the Laplacian term.
Test System Non-prob. VAE FM FM w/ symm. coupling
Two Delta Peaks 1.0* 0.25 0.091 -
@@
Buckling Beam - - 43.2 14.8
Allen-Cahn - - 255 244
*Theoretical result, a non-probabilistic model would always predict the mean 0.
Buckling:
To demonstrate the capability of flow matching to capture bifurcations with multiple solutions in a more realistic scenario, we test it on a buckling beam problem in 2D. We use an approach similar to brinke2025flow, which uses a random walk prior, an EGNN satorras2021n for interactions between nodes and a UNet for the time evolution. The symmetric coupling in this case consists of picking the closest reflection of the target trajectory. The results in Table 1 show that, as expected, symmetric coupling improves the results. See Figure 1(b,c) for an example of the predictions of a trained flow matching model compared to the two possible ground truths.
Allen-Cahn:
Finally, we test flow matching on the 1D Allen-Cahn equation with periodic boundary conditions, which shows complex bifurcation behavior, and in which the number of solutions can quickly become infinite, including (but not limited to) all possible circular shifts of a solution. The results in Table 1 show that using symmetric coupling to find the closest circular shift (determined using FFT cross-correlation), as well as the best reflection and sign flip, improves the results. The trained model additionally manages to recreate the pitchfork bifurcation diagram as the parameter $\mu$ is varied, see the appendix C for more detailed results of this system.
@@ 3 Roads:
This test system models bifurcation in a situation where there are two entities (can be nodes in a graph or something else) that need to coordinate. One can imagine that these are two people that are in a store with very narrow aisles. Both of them are standing right in front of an obstacle separating the aisles, and want to avoid this obstacle by either swerving left or right. However, they don’t want to bump into each other, so if one of them chooses the middle aisle (e.g., going left for the right person, going right for the left person), the other will not pick the same aisle. See Figure 2 for an illustration. This is important in prediction of, e.g., pedestrian dynamics, where 1) people coordinate to avoid bumping into each other, and 2) the distributions of their possible paths are highly multimodal when obstacles are involved; there are multiple ways to go around an obstacle, but none to go through it minartz2025discovering, brinke2025flow. We choose to model this as a very simple system with two input features and two output features. The input and output features $x_{i}^{(1)},x_{i}^{(2)}]$ and $y_{i}^{(1)},y_{i}^{(2)}]$ are two numbers that indicate the horizontal position of both entities before and after deciding their path, respectively. The possible values of $y_{i}^{(1)},y_{i}^{(2)}]$ are
@@ 4 Node Graph:
For this problem, the input is a graph consisting of 4 nodes, connected in a square. Each node has the same value $x_{i}\in[-100,100]$ as a node embedding. The output is an embedding for each node which is equal to $x_{i}\pm 5$, with connected nodes choosing the opposite sign. See Figure 3 for a visualization. This system is equivariant with respect to permutation. Additionally, the input is self-similar under some permutations, whereas each individual output is self-similar under fewer permutations, which means there is symmetry-breaking. Our dataset contains 2000 data points, of which 70% is used for training, 30% for testing. We sample $x_{i}$ uniformly from the range [50,50][-50,50].
@@
For this problem, we consider a beam that is attached to a base on the bottom, and then prescribe a vertical displacement to the top of the beam, while leaving the horizontal displacement free. For a small displacement, the beam will compress but stay straight, but when the displacement reaches a critical value, the system becomes unstable and the beam will buckle, which will make it bend to either the left or right side, with either option equally likely. We model the beam as a series of $n$ connected segments, which connect $n+1$ nodes, where node 0 is attached to the base. We consider $n$ ranging from 2 to 10. Each segment has a length $L_{i}$ and a stiffness $C_{i}$. At each node except for the top one the beam can bend, and therefore has a rotational stiffness $K_{i}$. The input $d$ is the vertical displacement applied to the top node. We solve the deformation of the beam by finding the Hencky strains $\epsilon_{i}$ and angles $q_{i}$ that minimize the total energy:
@@
Allen-Cahn:
The Allen-Cahn equation is a reaction-diffusion equation and describes the time evolution of a binary mixture of phases, e.g., ice and water at zero degrees Celsius. It is given by
@@
where $u$ is the phase field variable, representing the concentration of one of the phases. $\epsilon$ is parameter that controls size the of the interfaces between phases, and $u^{3}-\mu u$ is the derivative of a double-well potential with minima at $\pm\sqrt{\mu}$ if $\mu>0$, for $\mu\leq 0$ there is only one minimum at 0. We solve this equation on a 1D domain of length 1.0 with periodic boundary conditions, from $t=0$ to $t=100$, using a finite difference scheme, with 200 spatial points and a time step of 0.1. Each trajectory starts from $u=0$ homogeneously, however, because this is an unstable solution when $\mu>0$, we add a small random perturbation sampled from $\mathcal{N}(0,0.001)$ to each spatial point. We generate 400 trajectories in total, of which 300 are used for training and 100 for testing. For each trajectory, $\epsilon$ is randomly sampled from a loguniform distribution between 0.001 and 0.1., and $\mu$ from $\mathcal{U}(-0.1,1.0)$.

### v3

@@ Abstract
Bifurcation phenomena in nonlinear dynamical systems often lead to multiple coexisting stable solutions, particularly in the presence of symmetry breaking. Deterministic machine learning models are unable to capture this multiplicity, averaging over solutions and failing to represent lower-symmetry outcomes. In this work, we formalize the use of generative AI, specifically flow matching, as a principled way to model the full probability distribution over bifurcation outcomes. Our approach builds on existing techniques by combining flow matching with equivariant architectures and an optimal-transport-based coupling mechanism. We generalize equivariant flow matching to a symmetric coupling strategy that aligns predicted and target outputs under group actions, allowing accurate learning in equivariant settings. We validate our approach on a range of systems, from simple conceptual systems to physical problems such as buckling beams and the Allen–Cahn equation. The results demonstrate that the approach accurately captures multimodal distributions and symmetry-breaking bifurcations. Moreover, our results demonstrate that flow matching significantly outperforms non-probabilistic and variational methods. This offers a principled and scalable solution for modeling multistability in high-dimensional systems.
@@ 1 Introduction
Machine learning enables efficient data-driven modeling of complex systems, especially when traditional methods impose high complexity or are incomplete Senior et al. (2020); Brunton et al. (2020); Pfaff et al. (2020); Hendriks et al. (2025); Minartz et al. (2025); Brinke et al. (2025). Many such systems exhibit bifurcations that mark sudden changes in system behavior due to a small change in a control parameter. Bifurcations are crucial in fields such as fluid dynamics Ruelle and Takens (1971), climate science Armstrong McKay et al. (2022); Willcock et al. (2023); Rietkerk et al. (2021); Steffen et al. (2018), biology Sadria and Bury (2024), and crowd dynamics Muramatsu et al. (1999); Gu et al. (2025); Warren and Wirth (2022), where predicting transitions is essential. Many bifurcating systems exhibit multistability, where multiple distinct stable states coexist under the same input parameters, and the number of these states can change as the parameters vary. This phenomenon often arises due to symmetry breaking, where a system’s state after the bifurcation possesses fewer symmetries than before. This results in a multiplicity of possible solutions, all of which are equally valid.
For example, a fluid flow at high Reynolds number can start oscillating, breaking time-translation symmetry Ruelle and Takens (1971). A symmetric structure under compression can buckle in multiple directions, breaking rotation and reflection symmetry Bažant and Cedolin (2003). An even mixture of phases can separate into distinct regions, breaking translational and reflection symmetry. Symmetry breaking is also central to magnetism, superconductivity, and particle physics Beekman et al. (2019); Strocchi (2005), as well as the formation of enantiomers Kondepudi and Asakura (2001). In mechanical metamaterials, buckling is a design feature that enables programmable shape transformations and tunable mechanical responses through controlled symmetry breaking Kang et al. (2013); Azulay and Combescure (2024).
@@
Simulating such systems with machine learning is nontrivial and remains an open problem (Zhang et al., 2025, §2.9.1). Deterministic machine learning models can only produce the average of multiple solutions, resulting in nonphysical predictions that do not represent any true solution. Geometric deep learning approaches that rely on equivariant models preserve the system’s symmetry but cannot select asymmetric outcomes, limiting their ability to capture bifurcations. Approaches such as canonicalization Lawrence et al. (2025), input perturbation Hendriks et al. (2025), model ensembles Zou et al. (2025), deterministic machine learning models combined with pseudo-arclength continuation Fabiani et al. (2025); Della Pia et al. (2025), and sampling group operations to artificially break symmetry Xie (2024); Lawrence et al. (2025) offer partial solutions but often lack generalizability or precision.
We propose modeling the full output distribution using generative modeling. Generative models aim to learn mappings from simple base distributions (e.g., Gaussian noise) to complex, high-dimensional target distributions. We use generative modeling to parameterize the distribution of possible bifurcation outcomes given the input parameters.
@@
However, the probability distributions we require are singular: the set of allowed values (the support) lies on a subspace that has lower dimensionality than the input space, e.g., multiple Dirac deltas in 1D, or a circle in 2D space. This means the probability mass must be concentrated as sharply as possible around the allowed values. When the target distribution is singular, the necessary mapping becomes highly nonlinear: nearby points in the source distribution may map to distant points in the target space. Capturing such mappings is challenging for neural networks, as it requires representing high-frequency functions Rahaman et al. (2019). This is a fundamental limitation of direct generative approaches, which must learn a single, highly nonlinear transformation. For example, variational autoencoders (VAE) tend to generate blurry images due to their limitations in capturing highly concentrated, multimodal distributions Larsen et al. (2016); Bousquet et al. (2017); Dosovitskiy and Brox (2016).
@@
In contrast, iterative methods such as diffusion models, denoising processes, and flow matching approximate this complex mapping as a sequence of small integration steps. This iterative structure distributes the required nonlinearity across many stages, making the learning problem more tractable, as each step only needs to model a small, smooth transformation. As a result, these methods are better suited for (i) modeling singular distributions, where the probability mass is concentrated on a low-dimensional manifold, and (ii) multimodal distributions, where the probability mass must split far apart Salmona et al. (2022).
@@
Therefore, to address the challenge of modeling multistability, we utilize flow matching Lipman et al. (2022) to model the full output distribution. We show that flow matching excels at parameterizing distributions whose probability mass is highly concentrated in a low-dimensional subspace (e.g., very close to two Dirac deltas). We further systematically integrate flow matching with equivariant modeling to tackle symmetry-breaking. Additionally, we develop ‘symmetric coupling’, an approach to determine the optimal training target, similar to minibatch optimal transport Tong et al. (2023) and equivariant flow matching Klein et al. (2023); Song et al. (2023), that exploits the self-similarity of the input. This approach substantially improves the training and the quality of generated flow paths. We demonstrate our method on abstract and physical systems, including coin flips, buckling beams, and phase separation via the Allen–Cahn equation. Our approach effectively captures bifurcations and multistability, with and without symmetry breaking.
@@ 2 Methods
If an input $x$ is invariant (i.e., self-similar) under a subset of the group actions (i.e., $g_{x}\cdot x=x$ for all $g_{x}\in G_{x}$, with $G_{x}\subseteq G$), then equivariance implies that output $y$ is also invariant (i.e., $g_{x}\cdot y=y$). However, this preservation of symmetry inherently prevents a model from expressing symmetry-breaking behavior. To capture bifurcations and transitions into asymmetric states, equivariant models must be adapted to allow for symmetry-breaking solutions, which means the traditional equivariance condition in Equation (1) cannot hold for individual outputs after the bifurcation.
@@
The equivariance in Equation (1) then still holds for the set of solutions $\{g_{x}\cdot y|g_{x}\in G_{x}\}$. This is similar to the definition of ‘relaxed equivariance’ by Kaba et al. Kaba et al. (2023); Kaba and Ravanbakhsh (2023) in that both account for input symmetries via stabilizer subgroups (i.e., groups $G_{x}$ whose actions leave the input $x$ invariant), but it differs by preserving equivariance over sets of outputs (i.e., orbits of solutions), whereas relaxed equivariance modifies the equivariance condition itself, by allowing a correction with an element from the input’s stabilizer subgroup.
As a probability distribution.
@@
To ensure that the predictions by the model satisfy Equation (2), we use (i) a model that is equivariant under $G$ to parametrize the probability distribution and (ii) a prior $p(y_{0})$ that is invariant under $G$. This ensures that the generated probability distribution respects the symmetry of the problem Minartz et al. (2023).
@@
To model this probability distribution, we use flow matching Lipman et al. (2022), which learns a vector field $u(y_{t},t,x)$ that transforms samples from an uninformed prior $p(y_{0})$ into samples $y_{1}$ from the target distribution $p(y|x)$ over a continuous pseudo-time variable $t\in[0,1]$. The vector field is learned by minimizing the expected squared error between the model prediction and the target vector field along linear interpolation paths between prior and target samples.
Symmetric coupling.
@@
This is a way of straightening the flow paths similar to minibatch optimal transport Tong et al. (2023), but instead of finding the optimal coupling between a minibatch of the prior samples and target samples, we find the optimal coupling between one prior sample $y_{0}$ and all symmetric equivalents $g_{x}\cdot y_{1},g_{x}\in G_{x}$ of its corresponding target sample $y_{1}$. Our approach is similar to equivariant flow matching as described by Klein Klein et al. (2023) and Song Song et al. (2023), except we only consider the group $G_{x}$ that leaves the input $x$ invariant. If this group contains permutations or rotations, we can therefore follow their lead and use the Hungarian algorithm for permutations or the Kabsch algorithm for rotations. Other options are checking all possible reflections of the target or using FFT cross-correlation to determine the optimal periodic translation of the target.
@@
Figure 1: (a) Self-similarity and equivariance illustrated with the buckling beam problem, in which the input is the state at the current time step $t$ and the target output is the next time step $t+1$. $d$ is the vertical downward displacement of the beam tip relative to the total length of the beam, which is increased over time. The relevant symmetry group $G$ here consists of only two elements: identity and reflection in the y-axis. In this case, the self-similarity of the beam before buckling happens to be described by the same group, i.e., here $G_{x}=G$. (b) The $x$-coordinates of each node over the course of a trajectory, showing 50 predictions of a trained flow matching model (dashed lines) compared to the two possible ground truths (solid lines). (c) The same 50 predictions compared to the two possible ground truths, showing the beam deformation.
@@ 3 Experiments
For the 2-Delta Peaks dataset and the 4-node graph, we additionally show how symmetric coupling further improves the results. For the 2-Delta Peaks, even though the model is not equivariant, we can still use symmetric coupling to improve the results, since the targets and the prior are symmetric to a sign flip. Symmetric coupling then amounts to picking -1 or +1 depending on which is closer to the prior sample. For the 4-node graph, symmetric coupling means picking the permutation of the target that has the smallest Euclidean distance squared to the noise sample. Only two permutations are relevant, because there are only two possible solutions; the identity permutation and the one that swaps node 0 with 2 and node 1 with 3.
Table 1: Performance of various approaches for the different test systems. All metrics except for the Allen–Cahn are the Wasserstein distance of a set of 100 predictions to the actual distribution of allowed outcomes for each test data point, averaged over all test data points. For the Allen–Cahn system, we use the residual of the governing equation as a performance metric, acknowledging its limitations due to sensitivity to small fluctuations in the prediction, particularly in the Laplacian term.
Test System Non-prob. VAE FM FM*
Two Delta Peaks 1.0** 0.25 0.091 0.0041
@@
Buckling Beam - - 23.1 9.6
Allen–Cahn - - 255 244
*With symmetric coupling.
**Theoretical result, a non-probabilistic model would always predict the mean 0.
Buckling.
To demonstrate the capability of flow matching to capture bifurcations with multiple solutions in a more realistic scenario, we test it on a buckling beam problem in 2D. We use an approach similar to Brinke et al. (2025), which uses a random walk prior, an EGNN Satorras et al. (2021) for interactions between nodes and a UNet for the time evolution. The symmetric coupling in this case consists of picking the closest reflection of the target trajectory. The results in Table 1 show that, as expected, symmetric coupling improves the results. See Figure 1(b,c) for an example of the predictions of a trained flow matching model compared to the two possible ground truths.
Allen–Cahn.
Finally, we test flow matching on the 1D Allen–Cahn equation with periodic boundary conditions, which shows complex bifurcation behavior, and in which the number of solutions can quickly become infinite, including (but not limited to) all possible circular shifts of a solution. The results in Table 1 show that using symmetric coupling to find the closest circular shift (determined using FFT cross-correlation), as well as the best reflection and sign flip, improves the results. The trained model additionally manages to recreate the pitchfork bifurcation diagram as the parameter $\mu$ is varied, see the appendix C for more detailed results of this system.
@@ 3 Roads:
This test system models bifurcation in a situation where there are two entities (can be nodes in a graph or something else) that need to coordinate. One can imagine that these are two people that are in a store with very narrow aisles. Both of them are standing right in front of an obstacle separating the aisles, and want to avoid this obstacle by either swerving left or right. However, they don’t want to bump into each other, so if one of them chooses the middle aisle (e.g., going left for the right person, going right for the left person), the other will not pick the same aisle. See Figure 2 for an illustration. This is important in prediction of, e.g., pedestrian dynamics, where 1) people coordinate to avoid bumping into each other, and 2) the distributions of their possible paths are highly multimodal when obstacles are involved; there are multiple ways to go around an obstacle, but none to go through it Minartz et al. (2025); Brinke et al. (2025). We choose to model this as a very simple system with two input features and two output features. The input and output features $x_{i}^{(1)},x_{i}^{(2)}]$ and $y_{i}^{(1)},y_{i}^{(2)}]$ are two numbers that indicate the horizontal position of both entities before and after deciding their path, respectively. The possible values of $y_{i}^{(1)},y_{i}^{(2)}]$ are
@@ 4 Node Graph:
For this problem, the input is a graph consisting of 4 nodes, connected in a square. Each node has the same value $x_{i}$, as a node embedding, which is sampled uniformly from the range [50,50][-50,50]. The output is an embedding for each node which is equal to $x_{i}\pm 5$, with connected nodes choosing the opposite sign. See Figure 3 for a visualization. This system is equivariant with respect to permutation. Additionally, the input is self-similar under some permutations, whereas each individual output is self-similar under fewer permutations, which means there is symmetry-breaking. Our dataset contains 2000 data points, of which 70% is used for training, 30% for testing.
@@
For this problem, we consider a beam that is attached to a base on the bottom, and then prescribe a vertical displacement to the top of the beam, while leaving the horizontal displacement free. For a small displacement, the beam will compress but stay straight, but when the displacement reaches a critical value, the system becomes unstable and the beam will buckle, which will make it bend to either the left or right side, with either option equally likely. We model the beam as a series of $n$ connected segments, which connect $n+1$ nodes, where node 0 is attached to the base. We consider $n$ ranging from 2 to 10. Each segment has a length $L_{i}$ and an axial stiffness $C_{i}$. Individual segments can rotate relative to each other, reflected by a rotational stiffness $K_{i}$. The input $d$ is the vertical displacement applied to the top node. We solve the deformation of the beam by finding the Hencky strains $\epsilon_{i}$ and angles $q_{i}$ that minimize the total energy:
@@
Allen–Cahn:
The Allen–Cahn equation is a reaction-diffusion equation and describes the time evolution of a binary mixture of phases, e.g., ice and water at zero degrees Celsius. It is given by
@@
where $u$ is the phase field variable, representing the concentration of one of the phases. $\epsilon$ is parameter that controls the size of the interfaces between phases, and $u^{3}-\mu u$ is the derivative of a double-well potential with minima at $\pm\sqrt{\mu}$ if $\mu>0$, for $\mu\leq 0$ there is only one minimum at 0. We solve this equation on a 1D domain of length 1.0 with periodic boundary conditions, from $t=0$ to $t=100$, using a finite difference scheme, with 200 spatial points and a time step of 0.1. Each trajectory starts from $u=0$ homogeneously, however, because this is an unstable solution when $\mu>0$, we add a small random perturbation sampled from $\mathcal{N}(0,0.001)$ to each spatial point. We generate 400 trajectories in total, of which 300 are used for training and 100 for testing. For each trajectory, $\epsilon$ is randomly sampled from a loguniform distribution between 0.001 and 0.1, and $\mu$ from $\mathcal{U}(-0.1,1.0)$.